Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Thursday, November 12, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 12, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/12/98 (Thursday)

Segment 1 (Emidio to ANF) Report for 11/12/98. Excavation and Shoring of pits at the Culverts on Peace Valley road, EMP 25.45. Clearing activities occurred at the skip between EMP 21.73 and EMP 21.52. Tie -in welds were made on the Tejon bench at EMP 18.50. Excavation of tie-in trench at EMP 18.15. A coating crew worked at EMP 17.63. Welds were made on a section of pipe near EMP 16.77. Trenching continued on Grapevine Road northward to ~EMP 15.88. Station work continued at both Grapevine and Emidio Stations. Pipe was welded and x-rayed near EMP 14.53. Trenching, shoring and dewatering of the trench took place along Highway 166 near EMP .55. Road work and repairs were made on Access road A, between EMP 4.56 and 3.06.

Segment 2 (ANF to Sylmar) Report for 11/12/98): 2C (Templin Hwy to Castaic): Shading took place at the testing cutout tie-in south of Templin Hwy, EMP 50 89. Materials, including blocks for retaining wall construction and straw bales for revegetation were moved from storage sites in the northern half of Segment 2C to locations near Castaic. Revegetation work concentrated on the steep slopes on either side of EMP 54.74. Shading, F/O conduit -laying, and backfill continued south of EMP 55.46. The Ozzie padding machine tipped over during shading operation. No injuries occurred. A resultant spill of diesel fuel was quickly contained by crews. Work continued on the anchor structure at Lake Hughes block valve, EMP 56. 62. 2B (Castaic to Sylmar): Boring continue at the Feedmill Rd directional drill. Welders also prepared six-inch casing for F/O conduit. At Calgrove Blvd, bending and lowering-in of pipe occurred. Trenching took place north of Sierra Hwy in CalTrans property. South of Sierra Hwy, welding, x-ray, and coating reached EMP 72.76. Asphalt and slurry were poured at Sunshine Cyn.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 12 November 1998: 3A: Backfill crews worked on sunken areas of the right of way, near Sonora Avenue and Fletcher Avenue. A crew hooked up solar power to the block valve in Taylor Yard, MP 96.00. Welding, x-ray, and coating continued their work on the pipe at Hollywood Way and San Fernando Road. Final work on the bore pit at Tuxford was done, with the bore crew scheduled to set up tomorrow. Three brackets were hung on the Tujunga Wash Bridge crossing. Pipeline markers continued to be installed. NIGHT crews set up for an hour before deciding work needed to be rescheduled due to the need for additional equipment. 3B: No work occurred. 3C: Backfill continued to work on both the right of way and the road way near Grapevine Creek and Sepulveda Boulevard. The bore crew packed up and moved to Tuxford Street. Minor work occurred on the California Aqueduct crossing.

Segment 5 (Lynwood to Long Beach) Report for 12 November 1998: Directional drill at Highway 91/Compton Creek continued with 36" reamer. Two cased bores are continuing at Wilmington Refinery.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 5 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page